OpenStudy (anonymous):

what is the x in -2(7x-4)+2x=-4(x+1)

OpenStudy (highlyeducated87):

x=3/2

OpenStudy (anonymous):

First distribute to get.. \[−14x+8+2x=−4x−4\] Now get x to one side..

OpenStudy (anonymous):

So you get \[-14x+2x+4x=-4-8\]\[-8x=-12\]\[x=\frac{12}{8}\]This could be further simplified like this..\[x=\frac{3}{2}\]